OpenStudy (sissyedgar):

Which ordered pairs are solutions to the inequality 2x+y>−4? Select EACH correct answer. (−3, 0) (4, −12) (5, −12) (0, 1) (−1, −1) ♥ Thank you! ♥

satellite73 (satellite73):

lets check the first one is \(2\times -3+0\) bigger than \(-4\)?

OpenStudy (sissyedgar):

2 x -3 + 0 = -6 So no it isn't.

satellite73 (satellite73):

ok so that is out is \[2\times 4+(-12)\]bigger than \(-4\)?

OpenStudy (sissyedgar):

2 x 4 + (-12) = -4 So it's equal to

satellite73 (satellite73):

yeah so they are trying to trick you that one is out too since it say \(>\) not \(\geq\)

satellite73 (satellite73):

how about the next one \[2\times 5+(-12)\]

OpenStudy (sissyedgar):

2 x 5 + (-12) = -2 Which is greater than.

satellite73 (satellite73):

ok so that one is in

OpenStudy (sissyedgar):

2 x 0 + 1 = = -1 Also greater than. ♥ 2 x (-1) + (-1) = -3 Which is greater than too.

satellite73 (satellite73):

yes

OpenStudy (sissyedgar):

Awesome! Thankz! ♥